OpenStudy (anonymous):

How many and of what type are the solutions to x2 + 9x + 10 = 0? a. No real solutions b. Two identical rational solutions c. Two different rational solutions d. Two irrational solutions

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

I would say D. https://www.desmos.com/calculator

OpenStudy (anonymous):

well there will be 2 solutions because this is a trinomial and it has the power of 2

OpenStudy (igreen):

^ Exactly

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thx

OpenStudy (unklerhaukus):

check the discriminant the discriminant of x^2 + 9x + 10 = 0 ∆ = (9)^2-4(1)(10)

OpenStudy (anonymous):

ok

OpenStudy (unklerhaukus):

is the discriminant less than, equal to or grater than zero?

OpenStudy (unklerhaukus):

if \(Δ=0\) there is two equal solutions if \(Δ>0\) there are two different solutions if \(Δ<0\) there are no real solutions if \(∆\neq0=d^2\) (for some integer \(d\)) the solutions are rational